# Pluto and Charon: A Time and Memory Efficient Collaborative Edge AI Framework for Personal LLMs Fine-Tuning

### v1

@@
Abstract.
Large language models (LLMs) have unlocked a plethora of powerful applications at the network edge, such as intelligent personal assistants. Data privacy and security concerns have prompted a shift towards edge-based fine-tuning of personal LLMs, away from cloud reliance. However, this raises issues of computational intensity and resource scarcity, hindering training efficiency and feasibility. While current studies investigate parameter-efficient fine-tuning (PEFT) techniques to mitigate resource constraints, our analysis indicates that these techniques are not sufficiently resource-efficient for edge devices. Other studies focus on exploiting the potential of edge devices through resource management optimization, yet are ultimately bottlenecked by the resource wall of individual devices.
To tackle these challenges, we propose Pluto and Charon (PAC), a time and memory efficient collaborative edge AI framework for personal LLMs fine-tuning. PAC breaks the resource wall of personal LLMs fine-tuning with a sophisticated algorithm-system co-design. (1) Algorithmically, PAC implements a personal LLMs fine-tuning technique that is efficient in terms of parameters, time, and memory. It utilizes Parallel Adapters to circumvent the need for a full backward pass through the LLM backbone. Additionally, an activation cache mechanism further streamlining the process by negating the necessity for repeated forward passes across multiple epochs. (2) Systematically, PAC leverages edge devices in close proximity, pooling them as a collective resource for in-situ personal LLMs fine-tuning, utilizing a hybrid data and pipeline parallelism to orchestrate distributed training. The use of the activation cache eliminates the need for forward pass through the LLM backbone, enabling exclusive fine-tuning of the Parallel Adapters using data parallelism. Extensive evaluation based on prototype implementation demonstrates that PAC remarkably outperforms state-of-the-art approaches, achieving up to 8.64×8.64$\times$8.64 × end-to-end speedup and up to $88.16\%$ reduction in memory footprint.
Edge intelligence, large language model, parameter-efficient fine-tuning, pipeline parallelism, data parallelism, parallel processing
$\bigstar$: Equal contributions. $\dagger$: Corresponding author.
journalyear: 2024copyright: rightsretainedconference: The 53rd International Conference on Parallel Processing; August 12–15, 2024; Gotland, Swedenbooktitle: The 53rd International Conference on Parallel Processing (ICPP ’24), August 12–15, 2024, Gotland, Swedendoi: 10.1145/3673038.3673043isbn: 979-8-4007-1793-2/24/08
1. Introduction
@@
Large language models (LLMs) (Vaswani et al., 2017; Raffel et al., 2020; Lewis et al., 2019) have ushered in a revolution in machine intelligence, owing to their exceptional capabilities in a wide range of machine learning tasks. While born on datacenter warehouse, LLMs have quickly sunk to edge devices and facilitated a range of intelligent applications at the network edge, such as intelligent personal assistants (IPAs) which are software agents that can augment individuals’ abilities, complete complicated tasks, and even satisfy emotional needs. A recent survey (Li et al., 2024) targeting LLM-based IPAs has revealed that over $80\%$ of industry experts believe that, owing to the sensitive and privacy-critical nature of user data, personal LLMs should be fully (or primarily) hosted at the edge in order to enable privacy-preserving model personalization and serving. Figure 1 illustrates the scenario of hosting a personal LLM-based intelligent agent within a smart home. A personal LLM agent provides users with high-performance, privacy-preserving intelligent services. Meanwhile, the agent also tracks user interactions, learns from experiences, and extracts knowledge to fine-tune the personal LLMs and further enhance the service quality.
@@
Figure 1. An illustration of hosting personal LLM-based intelligent agents within a smart home.
@@
While the serving of LLMs on edge devices has been made feasible through careful engineering (Guo et al., 2023; Xu et al., 2023; Ye et al., 2024a), fine-tuning these models remains significantly challenging due to the resource-intensive nature of LLM training. Towards alleviating the resource challenges, some research works (Cai et al., 2023; Miao et al., 2024) have explored parameter-efficient fine-tuning (PEFT) techniques, such as Adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021), which modify less than $2\%$ of the model’s parameters, thereby reducing resource requirements. Although these techniques are highly parameter-efficient, our analysis observes that they are not resource-efficient enough for edge environments. The inefficiency stems from Adapters and LoRA embedding trainable structures in the LLM backbone, necessitating complete backward passes through the LLMs during backpropagation. As we will empirically show in §2, fine-tuning a popular LLM of T5-Base (0.25B) by Google with PEFT techniques can only reduce computational overhead by up to $30\%$ compared to full model fine-tuning. In practice, fine-tuning the T5-Base on a typical edge device (e.g., NVIDIA Jetson Nano (jet, 2019)) still demands a minimum of $72.6$ minutes per training epoch employing LoRA. Moreover, on-device fine-tuning is severely hindered by the memory wall of a single device. Predominant techniques require substantial memory to accommodate both model parameters and intermediate results. The observed memory expense for fine-tuning LLMs like T5-Large (0.74B), which exceeds $7.1$ GB with LoRA and $6.8$ GB with Adapters, is often unaffordable as typical mobile devices only possess 4-12GB DRAMs in total to run both system software and applications.
@@
Other leading researchers have explored designing sophisticated resource management mechanisms (e.g., CPU-DSP co-execution (Xu et al., 2022), memory budget adapting (Gim and Ko, 2022; Wang et al., 2022)) to leverage native resources, but are still bottlenecked by the intrinsic resource shortage of single device. To break the resource wall of a single device, we alternatively observe that prevalent edge environments like smart homes usually comprise a group of trusted idle devices beyond a single terminal (e.g., phones and smart-home devices). These accompanying devices are typically in physical proximity and can be associated as a resource augmentation for in-situ personal LLMs fine-tuning.
@@
As motivated, in this paper, we introduce Pluto and Charon (PAC), a time and memory efficient collaborative edge AI framework for personal LLMs fine-tuning. PAC’s contribution goes beyond merely leveraging distributed edge devices, instead it breaks the resource wall of in-situ personal LLMs fine-tuning with a sophisticated algorithm-system co-design:
@@
$\bullet$ (Algorithm) We evaluate two predominant PEFT techniques, Adapters and LoRA, and reveal that although parameter efficient, these techniques do not achieve sufficient resource efficiency. In light of the side-tuning (Zhang et al., 2020) techniques, we employ not only parameter but also time and memory-efficient personal LLMs fine-tuning techniques with Parallel Adapters, which provides a dedicated gradient "highway" for the trainable parameters. Additionally, our Parallel Adapters stand out from other PEFT techniques by preserving the invariant intermediate activations from the LLM backbone for any given input sequence. By reusing these cached activations across multiple epochs, PAC increases resource efficiency and reduces fine-tuning latency by eliminating repetitive forward propagation through the LLM backbone.
@@
$\bullet$ (System) We leverage edge devices in physical proximity and associate them as an edge resource pool for in-situ personal LLMs fine-tuning. Our fine-tuning process can be divided into two phases: (1) For the first epoch, the LLMs backbone, augmented with Parallel Adapters, is fine-tuned across multiple edge devices. To enhance scalability and training throughput, a hybrid parallelism approach that combines the merits of both data and pipeline parallelism is employed by PAC as a principle to manage collaborative training across multiple edge devices. (2) In subsequent fine-tuning epochs, the activation cache obviates the need for forward propagation through the LLM backbone, allowing for the exclusive fine-tuning of our Parallel Adapters using data parallelism.
@@
We implement PAC in realistic testbeds with a cluster of edge devices. Extensive evaluations across three LLMs demonstrate that PAC not only accelerates fine-tuning up to 8.64×8.64$\times$8.64 × faster than existing state-of-the-art methods but also significantly lowers the peak memory footprint by up to $88.13\%$, without sacrificing model performance. The main contributions are summarized as follows.
  • We carry out extensive measurement studies on predominant PEFT techniques on resource-constrained edge devices and demonstrate that they are not sufficiently resource-efficient.
@@
  • We design a not only parameter but also resource efficient LLM fine-tuning technique for resource-limited edge environments.
  • We propose a time and memory efficient collaborative edge AI framework PAC for the in-situ fine-tuning of personal LLMs, which combines sophisticated algorithm-system co-design.
@@
  • We implement PAC and evaluate it in realistic edge testbeds. Experimental results show up to 8.64×8.64$\times$8.64 × fine-tuning acceleration and $88.16\%$ memory reduction without sacrificing performance compared to state-of-the-art methods.
Figure 2. Illustration of the model structures with two PEFT.
2. MOTIVATION AND PRELIMINARIES
2.1. Transformer-Based LLMs and Fine-Tuning
[…]
@@ 2.1. Transformer-Based LLMs and Fine-Tuning
Personal LLMs Fine-Tuning. The training of LLMs typically consists of two stages: pre-training and fine-tuning. Before being deployed for specific tasks, language models are often pre-trained on extensive text datasets containing vast linguistic data. The pre-training process enables the model to acquire a general understanding of linguistic structure and patterns that are widely applicable. The fine-tuning adapts the pre-trained model to various, concrete downstream language tasks such as intelligent personal assistants. During actual deployment, the data required for fine-tuning is often generated at the user end, which can carry significant concerns regarding data security and privacy. In recent years, in-situ learning on edge devices (Patil et al., 2022; Lin et al., 2022; Gim and Ko, 2022; Wang et al., 2022) has emerged as a promising approach for customizing LLMs while preserving user data fully in-situ.
@@
Figure 3. The comparison of floating point of operations (FLOPs). Mini-batch size: 16; sequence length: 128.
@@
Table 1. The breakdown of memory footprint. "Activations" contain the intermediate results and optimizer states. Model: T5-Large; mini-batch size: 16; sequence length: 128.
@@
Full model fine-tuning updates all parameters of an LLM for a specific downstream task. However, it is impractical for adapting an LLM to multiple distinct downstream tasks, as each target task would require maintaining a separate LLM with whole parameters. Some leading researchers have proposed parameter-efficient fine-tuning (PEFT) techniques (Lester et al., 2021; Houlsby et al., 2019; Hu et al., 2021; Liu et al., 2024) which adapt a small subset of the LLM parameters or a set of newly added parameters for each new task. Adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021) are two of the most widely used PEFT techniques. Figure 2 illustrates how the transformer layer structure incorporates these two techniques. Specifically, adapters are compact bottleneck modules inserted at the end of each transformer layer. Similarly, LoRA injects trainable low-rank matrices into a frozen pre-trained model. These decompose the weight matrix parameter updates into two learnable low-rank matrices. Extensive experiments have demonstrated that these PEFT techniques can achieve performance comparable to full fine-tuning.
@@
Although these PEFT techniques can greatly reduce the number of trainable parameters (around $98\%$), our analysis has revealed that they do not significantly decrease the computational and memory requirements during training. Figure 3 illustrates the floating point of operations (FLOPs) of different fine-tuning techniques and inference. Adapters and LoRA exhibit a limited reduction in computation (around $30\%$). Table 1 summarizes the memory footprint breakdown for T5-Large. Although Adapters and LoRA minimize the gradient memory footprint by restricting the number of trainable parameters, the memory consumed by activations still constitutes substantial overhead, with a maximum reduction of only $36\%$. The reason is that both Adapters and LoRA introduce trainable structures within the LLM backbone, such as at the end of each transformer block or as bypasses to linear layers. Computing gradients for trainable parameters via backpropagation involves traversing the LLM backbone, compromising the efficiency of PEFT techniques due to the additional computational overhead and memory required to maintain considerable intermediate activations in LLM backbone.
@@
2.2. Personal LLMs Fine-Tuning with Resource-Constrained Edge Devices
@@
On-device fine-tuning enables leveraging idle resources at the edge while fully preserving user data privacy (Patil et al., 2022; Lin et al., 2022; Gim and Ko, 2022; Wang et al., 2022). This paradigm is widely adopted in privacy-sensitive edge computing applications. However, the resource-intensive nature of LLMs fine-tuning presents two significant challenges for resource-limited edge devices: (1) The computational capabilities of edge devices are constrained. Edge devices often face stark computational constraints compared to the powerful accelerators available in cloud datacenters. The Jetson Nano (jet, 2019), a specialized platform for edge AI, peaks at a mere $0.47$ TFLOPS, a tiny fraction of the $312$ TFLOPS achievable with NVIDIA’s A100 GPU typically found in data centers. Fine-tuning a T5-Base model with Adapters on a single Jetson Nano requires an epoch time of 72.6 minutes, which is 175.5×175.5$\times$175.5 × longer than that in a NVIDIA A100 GPU, showing the fundamental contradiction between intensive LLM fine-tuning workload and constrained on-board resources. (2) On-device fine-tuning is hindered by the memory wall. As shown in Table 1, fine-tuning the T5-Large model incurs a peak memory footprint that is often unaffordable for edge devices. For instance, although PEFT techniques such as Adapters and LoRA adjust only approximately $2\%$ of the parameters, they still require substantial memory $6.89$ GB and $7.13$ GB respectively. Compared to full model fine-tuning, which requires over $10$ GB, these techniques reduce memory usage by only $36\%$, often insufficient for typical mobile devices with 4-12 GB DRAM to run system software and applications.
@@
To break the resource wall of a single edge device, in our work, we alternatively observe that prevalent edge scenarios usually comprise a group of trusted idle devices beyond a single terminal. These accompanying devices are typically located in close physical proximity, such as being connected to the same local area network (LAN), and can be utilized as a resource augmentation for in-situ LLMs fine-tuning acceleration. While several pioneering research works (Ye et al., 2024a; Wei et al., 2024) have delved into collaborative edge computing to overcome resource limitations faced by edge devices, the majority of these works primarily focus on LLMs inference. Other studies (Cai et al., 2023; Xu et al., 2024) employing federated learning for fine-tuning LLMs with collaborative edge devices primarily address the dissolution of data silos, rather than resource augmentation within LANs.
Figure 4. PAC workflow.
3. System Overview

### v2

# Resource-Efficient Personal Large Language Models Fine-Tuning with Collaborative Edge Computing

@@
Abstract
Large language models (LLMs) have enabled transformative applications at the network edge, such as intelligent personal assistants. However, data privacy and security concerns necessitate a shift from cloud-centric paradigms to edge-based finetuning for personal LLMs. This transition is significantly hindered by intensive computational requirements and inherent resource scarcity, creating a "resource wall" that compromises training efficiency and feasibility. While current parameter-efficient finetuning (PEFT) and resource management strategies attempt to mitigate these constraints, they remain insufficient for the limited capacities of individual edge devices. To address these challenges, we propose PAC+, a resourceefficient collaborative edge AI framework for in-situ personal LLM fine-tuning. PAC+ overcomes the resource bottlenecks through a sophisticated algorithm-system codesign: (1) Algorithmically, PAC+ introduces a fine-tuning technique optimized for parameters, time, and memory. It utilizes Parallel Adapters to circumvent the need for a full backward pass through the LLM backbone. Furthermore, an activation cache mechanism streamlines the process by negating redundant forward passes across multiple epochs. (2) Systematically, PAC+ aggregates proximate edge devices into a collective resource pool, employing hybrid data and pipeline parallelism to orchestrate distributed training. By leveraging the activation cache, PAC+ enables the exclusive fine-tuning of Parallel Adapters via data parallelism, effectively bypassing the backbone’s constraints. Extensive evaluation of the prototype implementation demonstrates that PAC+ significantly outperforms existing collaborative edge training systems, achieving up to a $9.7\times$ end-to-end speedup. Furthermore, compared to mainstream LLM fine-tuning algorithms, PAC+ reduces memory footprint by up to $88.16\%$.
I Introduction
@@
Large language models (LLMs) [39, 36, 22] have ushered in a revolution in machine intelligence, owing to their exceptional capabilities in a wide range of machine learning tasks. While born on datacenter warehouse, LLMs have quickly sunk to edge devices and facilitated a range of intelligent applications at the network edge, such as intelligent personal assistants (IPAs) which are software agents that can augment individuals’ abilities, complete complicated tasks, and even satisfy emotional needs. A recent survey [25] targeting LLM-based IPAs has revealed that over $80\%$ of industry experts believe that, owing to the sensitive and privacy-critical nature of user data, personal LLMs should be fully (or primarily) hosted at the edge in order to enable privacy-preserving model personalization and serving. Figure 1 illustrates the scenario of hosting a personal LLM-based intelligent agent within a smart home. A personal LLM agent provides users with high-performance, privacy-preserving intelligent services. Meanwhile, the agent also tracks user interactions, learns from experiences, and extracts knowledge to fine-tune the personal LLMs and further enhance the service quality.
@@
Figure 1: An illustration of hosting personal LLM-based intelligent agents within a smart home.
@@
While the serving of LLMs on edge devices has been made feasible through careful engineering [9, 43, 45], fine-tuning these models remains significantly challenging due to the resource-intensive nature of LLM training. Towards alleviating the resource challenges, some research works [2, 29] have explored parameter-efficient fine-tuning (PEFT) techniques, such as Adapters [13] and LoRA [16], which modify less than $2\%$ of the model’s parameters, thereby reducing resource requirements. Although these techniques are highly parameter-efficient, our analysis observes that they are not resource-efficient enough for edge environments. The inefficiency stems from Adapters and LoRA embedding trainable structures in the LLM backbone, necessitating complete backward passes through the LLMs during backpropagation. As we will empirically show in §II, fine-tuning a popular LLM of T5-Base (0.25B) by Google with PEFT techniques can only reduce computational overhead by up to $30\%$ compared to full model fine-tuning. In practice, fine-tuning the T5-Base on a typical edge device (e.g., NVIDIA Jetson Nano [18]) still demands a minimum of $72.6$ minutes per training epoch employing LoRA. Moreover, on-device fine-tuning is severely hindered by the memory wall of a single device. Predominant techniques require substantial memory to accommodate both model parameters and intermediate results. The observed memory expense for fine-tuning LLMs like T5-Large (0.74B), which exceeds $7.1$ GB with LoRA and $6.8$ GB with Adapters, is often unaffordable as typical mobile devices only possess 4-12GB DRAMs in total to run both system software and applications.
@@
Other leading researchers have explored designing sophisticated resource management mechanisms (e.g., CPU-DSP co-execution [42], memory budget adapting [7, 40]) to leverage native resources, but are still bottlenecked by the intrinsic resource shortage of single device. To break the resource wall of a single device, we alternatively observe that prevalent edge environments like smart homes usually comprise a group of trusted idle devices beyond a single terminal (e.g., phones and smart-home devices). These accompanying devices are typically in physical proximity and can be associated as a resource augmentation for in-situ personal LLMs fine-tuning.
@@
As motivated, in this paper, we introduce PAC+, a resource-efficient collaborative edge AI framework for personal LLMs fine-tuning. PAC+’s contribution goes beyond merely leveraging distributed edge devices, instead it breaks the resource wall of in-situ personal LLMs fine-tuning with a sophisticated algorithm-system co-design:
@@
$\bullet$ (Algorithm) We evaluate two predominant PEFT techniques, Adapters and LoRA, and reveal that although parameter efficient, these techniques do not achieve sufficient resource efficiency. In light of the side-tuning [51] techniques, PAC+ leverages the existing Parallel Adapters approach [10, 37] to realize not only parameter but also time and memory-efficient personal LLMs fine-tuning. This technique provides a dedicated gradient "highway" for the trainable parameters, significantly reducing the computational and memory overhead of backpropagation on the LLM backbone. By employing low-precision quantization on the LLM backbone and tailored initialization methods for the Parallel Adapters, PAC+ further minimize the memory and time costs of LLM fine-tuning. Additionally, our Parallel Adapters stand out from other PEFT techniques by preserving the invariant intermediate activations from the LLM backbone for any given input sequence. By reusing these cached activations across multiple epochs, PAC+ increases resource efficiency and reduces fine-tuning latency by eliminating repetitive forward propagation through the LLM backbone.
@@
$\bullet$ (System) PAC+ leverage edge devices in physical proximity and associate them as an edge resource pool for in-situ personal LLMs fine-tuning. Our fine-tuning process can be divided into two phases: (1) For the first epoch, the LLMs backbone, augmented with Parallel Adapters, is fine-tuned across multiple edge devices. To enhance scalability and training throughput, a hybrid parallelism approach that combines the merits of both data and pipeline parallelism is employed by PAC+ as a principle to manage collaborative training across multiple edge devices. A novel heterogeneity-aware planning algorithm is introduced for multidimensional resources optimization. (2) In subsequent fine-tuning epochs, the activation cache obviates the need for forward propagation through the LLM backbone, allowing for the exclusive fine-tuning of our Parallel Adapters using data parallelism.
@@
We implement PAC+ in realistic testbeds with a cluster of edge devices. Extensive evaluations across three LLMs demonstrate that PAC+ not only accelerates fine-tuning, achieving up to a $9.7\times$ speedup over existing collaborative edge training systems, but also reduces memory footprint by up to $88.16\%$ compared to mainstream LLM fine-tuning algorithms.
The main contributions are summarized as follows.
@@
  • We conduct extensive measurement studies on prevalent PEFT techniques, demonstrating their inefficiency, and design a not only parameter but also time and memory efficient LLM fine-tuning technique for resource-limited edge environments.
  • We analyze and design low-precision quantization for the LLM backbone and tailored initialization methods for the Parallel Adapters, enabling further optimization of both memory and time overhead in LLM fine-tuning.
  • We design a hybrid parallel fine-tuning architecture to leverage edge devices in physical proximity, complemented by a novel heterogeneity-aware planning algorithm for optimizing multidimensional resources.
  • We design a system cache mechanism for invariant intermediate activations, eliminating the need for a forward pass through the LLM backbone and enabling exclusive fine-tuning of the Parallel Adapters using data parallelism.
@@
  • We implement PAC+ and evaluate it in realistic edge testbeds. Experimental results show up to $9.7\times$ fine-tuning acceleration and $88.16\%$ memory reduction without sacrificing performance compared to state-of-the-art methods.
Figure 2: Illustration of the model structures with two PEFT.
II MOTIVATION AND PRELIMINARIES
II-A Transformer-Based LLMs and Fine-Tuning
Transformer-Based LLMs. Transformer-based LLMs have gained prominence in various language-related applications due to their impressive performance. These models consist of multiple Transformer layers, each comprising two main components: the Multi-head Attention and the Feed Forward block. The Multi-head Attention block utilizes linear layers to generate query (Q), key (K), and value (V) matrices for each attention head, allowing for independent self-attention computations. The outputs of these attention heads are then concatenated and processed through a final linear layer. The Feed Forward block involves two linear operations that expand the hidden dimension size then reduce it back.
@@ II-A Transformer-Based LLMs and Fine-Tuning
Personal LLMs Fine-Tuning. The training of LLMs typically consists of two stages: pre-training and fine-tuning. Before being deployed for specific tasks, language models are often pre-trained on extensive text datasets containing vast linguistic data. The pre-training process enables the model to acquire a general understanding of linguistic structure and patterns that are widely applicable. The fine-tuning adapts the pre-trained model to various, concrete downstream language tasks such as intelligent personal assistants. During actual deployment, the data required for fine-tuning is often generated at the user end, which can carry significant concerns regarding data security and privacy. In recent years, in-situ learning on edge devices [34, 26, 7, 40] has emerged as a promising approach for customizing LLMs while preserving user data fully in-situ.
@@
Figure 3: The comparison of floating point of operations (FLOPs). Mini-batch size: 16; sequence length: 128.
TABLE I: The breakdown of memory footprint. "Activations" contain the intermediate results and optimizer states. Model: T5-Large; mini-batch size: 16; sequence length: 128.
@@
Full model fine-tuning updates all parameters of an LLM for a specific downstream task. However, it is impractical for adapting an LLM to multiple distinct downstream tasks, as each target task would require maintaining a separate LLM with whole parameters. Some leading researchers have proposed parameter-efficient fine-tuning (PEFT) techniques [21, 13, 16, 27] which adapt a small subset of the LLM parameters or a set of newly added parameters for each new task. Adapters [13] and LoRA [16] are two of the most widely used PEFT techniques. Figure 2 illustrates how the transformer layer structure incorporates these two techniques. Specifically, adapters are compact bottleneck modules inserted at the end of each transformer layer. Similarly, LoRA injects trainable low-rank matrices into a frozen pre-trained model. These decompose the weight matrix parameter updates into two learnable low-rank matrices. Extensive experiments have demonstrated that these PEFT techniques can achieve performance comparable to full fine-tuning.
@@
Although these PEFT techniques can greatly reduce the number of trainable parameters (around $98\%$), our analysis has revealed that they do not significantly decrease the computational and memory requirements during training. Figure 3 illustrates the floating point of operations (FLOPs) of different fine-tuning techniques and inference. Adapters and LoRA exhibit a limited reduction in computation (around $30\%$). Table I summarizes the memory footprint breakdown for T5-Large. Although Adapters and LoRA minimize the gradient memory footprint by restricting the number of trainable parameters, the memory consumed by activations still constitutes substantial overhead, with a maximum reduction of only $36\%$. The reason is that both Adapters and LoRA introduce trainable structures within the LLM backbone, such as at the end of each transformer block or as bypasses to linear layers. Computing gradients for trainable parameters via backpropagation involves traversing the LLM backbone, compromising the efficiency of PEFT techniques due to the additional computational overhead and memory required to maintain considerable intermediate activations in LLM backbone.
Figure 4: PAC+ workflow.
@@
II-B Personal LLMs Fine-Tuning with Resource-Constrained Edge Devices
@@
On-device fine-tuning enables leveraging idle resources at the edge while fully preserving user data privacy [34, 26, 7, 40]. This paradigm is widely adopted in privacy-sensitive edge computing applications. However, the resource-intensive nature of LLMs fine-tuning presents two significant challenges for resource-limited edge devices: (1) The computational capabilities of edge devices are constrained. Edge devices often face stark computational constraints compared to the powerful accelerators available in cloud datacenters. The Jetson Nano [18], a specialized platform for edge AI, peaks at a mere $0.47$ TFLOPS, a tiny fraction of the $312$ TFLOPS achievable with NVIDIA’s A100 GPU typically found in data centers. Fine-tuning a T5-Base model with Adapters on a single Jetson Nano requires an epoch time of 72.6 minutes, which is $175.5\times$ longer than that in a NVIDIA A100 GPU, showing the fundamental contradiction between intensive LLM fine-tuning workload and constrained on-board resources. (2) On-device fine-tuning is hindered by the memory wall. As shown in Table I, fine-tuning the T5-Large model incurs a peak memory footprint that is often unaffordable for edge devices. For instance, although PEFT techniques such as Adapters and LoRA adjust only approximately $2\%$ of the parameters, they still require substantial memory $6.89$ GB and $7.13$ GB respectively. Compared to full model fine-tuning, which requires over $10$ GB, these techniques reduce memory usage by only $36\%$, often insufficient for typical mobile devices with 4-12 GB DRAM to run system software and applications.
@@
To break the resource wall of a single edge device, in our work, we alternatively observe that prevalent edge scenarios usually comprise a group of trusted idle devices beyond a single terminal. These accompanying devices are typically located in close physical proximity, such as being connected to the same local area network (LAN), and can be utilized as a resource augmentation for in-situ LLMs fine-tuning acceleration. While several pioneering research works [45, 41] have delved into collaborative edge computing to overcome resource limitations faced by edge devices, the majority of these works primarily focus on LLMs inference. Other studies [2, 44] employing federated learning for fine-tuning LLMs with collaborative edge devices primarily address the dissolution of data silos, rather than resource augmentation within LANs.
II-C Technical Challenges
According to the aforementioned analysis and discussion, we can summarize two key challenges inherent to fine-tuning personal LLMs on resource-constrained edge devices:
(1) How to boost the time and memory efficiency for edge fine-tuning? The current predominant PEFT techniques like LoRA and Adapters, while parameter-efficient, are not inherently resource-efficient. These approaches still require substantial computational and memory resources for deployment. This poses a fundamental contradiction, as edge devices are typically constrained by cost and power considerations, and thus possess highly limited computational and memory capabilities. Therefore, we need to exploit a PEFT technique that is inherently more time and memory efficient, and well-suited for edge environments.
(2) How to orchestrate resource-efficient collaborative computing across multiple edge devices? Leveraging multiple edge devices for distributed fine-tuning is not trivial. Careful selection of parallel architectures and parallelism planning strategies is necessary to maximize resource utilization and avoid out-of-memory issues. Furthermore, an algorithm-system co-design approach is required to seamlessly integrate the resource-efficient PEFT techniques with our collaborative edge system. This holistic design is crucial for enabling scalable and high-performance personal LLMs fine-tuning in resource-constrained edge environments.
III System Overview